### v2

# RomanSetu: Efficiently unlocking multilingual capabilities of Large Language Models via Romanization

Abstract
This study addresses the challenge of extending Large Language Models (LLMs) to non-English languages using non-Roman scripts. We propose an approach that utilizes the romanized form of text as an interface for LLMs, hypothesizing that its frequent informal use and shared tokens with English enhance cross-lingual alignment. Our approach involves the continual pretraining of an English LLM like Llama 2 (Touvron et al., 2023) on romanized text of non-English, non-Roman script languages, followed by instruction tuning on romanized data. The results indicate that romanized text not only reduces token fertility by 2x-4x but also matches or outperforms native script representation across various NLU, NLG, and MT tasks. Moreover, the embeddings computed on romanized text exhibit closer alignment with their English translations than those from the native script. Our approach presents a promising direction for leveraging the power of English LLMs in languages traditionally underrepresented in NLP.
RomanSetu: Efficiently unlocking multilingual capabilities of
Large Language Models via Romanization
[…]
1 Introduction
Large language models (LLMs) demonstrate remarkable proficiency across a broad spectrum of natural language processing (NLP) tasks, as evidenced by various studies (Liu et al., 2023; Chung et al., 2022; Chowdhery et al., 2022; Wei et al., 2022; Goyal et al., 2022b). They excel not only in tasks for which they were explicitly trained but also in those they were not trained for. This achievement is mainly due to the availability of corpora (Wenzek et al., 2019; Abadji et al., 2021; Suárez et al., 2019) as well as the advancements in LLMs that leverage these datasets for pretraining (Touvron et al., 2023; Workshop et al., 2022; Chowdhery et al., 2022). Despite their proficiency in English, these models typically demonstrate reduced effectiveness when applied to non-English languages, highlighting a significant challenge in extending their benefits to non-English languages.
Figure 1: Performance of the BaseLLM (Llama 2 Base), our continually pretrained model (CPT), and our instruction finetuned model (IFT), in both native (N) and romanized (R) script settings. The average scores for different tasks (left) and languages (right) are compared in each radar chart. For CPT, we show 3-shot results where available, else 1-shot results. For IFT, we use zero-shot results.
The English-heavy LLMs (Touvron et al., 2023; Jiang et al., 2023; Zhang et al., 2022) still have some representation from other languages due to data leakage while creating the pre-training dataset, particularly for languages which use the same script as English i.e., the Roman script. This script sharing enables cross-lingual transfer and bestows some of these LLM capabilities to these languages. For languages using non-Latin scripts, the data representation is very limited to non-existent. Tokenization of text in the languages exhibits high fertility (Ács, 2020) and byte-level representation (Artetxe et al., 2019). Hence, these LLMs perform poorly on most of these languages, and inefficient tokenization also leads to high inference latency as well as inability to process long sequences. This disparity in performance raises a critical question: how can we extend the capabilities of LLMs to the languages written in non-Latin scripts?
A widely explored solution is extension of the tokenizer vocabulary to incorporate new languages and continual pre-training on native language data (Cui et al., 2023; Nguyen et al., 2023; Minixhofer et al., 2021). This approach is computationally demanding, since the models need to be continually pre-trained long enough to effectively learn the new embeddings and align the representations of English and the new language. Furthermore, this approach requires the availability of large volumes of text corpora.
In this work, we explore an alternative approach for more efficient knowledge transfer. Instead of utilizing the native script, we use the romanized form of text as the interface to the LLMs in our approach RomanSetu. [^1] The word setu in Sanskrit means bridge, referring to the Roman script serving as a bridge between English and other languages. The adoption of romanized representation is justified for several reasons. In many languages, romanized text has been frequently used in informal settings and on social media in recent history. This usage creates the potential for the inclusion of some romanized data representation in the pre-training corpus. Additionally, code-mixing with English is a common occurrence, and the romanized form shares tokens with English. This leads us to hypothesize that the romanized form is better aligned with English than the native script, thereby facilitating a more effective transfer from English.
In our approach, we first continually pre-train a English LLM like Llama 2 (Touvron et al., 2023) on romanized text for a language generated via transliteration. Subsequently, we perform instruction tuning on this continually pretrained model. We experiment with multiple languages from two language families and evaluate our approach on multiple NLU, NLG and MT benchmarks.
Our experiments and analysis indicate that:
  • The fertility of romanized text is 2x-4x smaller than native text, making the romanized form far more efficient than the native script.
  • The embeddings of the romanized text are closer to the embeddings of their English translations, compared to the embeddings of their native script equivalents, suggesting the former’s suitability for cross-lingual transfer.
  • Results across NLU, NLG, and MT tasks shows that multilingual instruction finetuning on romanized data produces results comparable or better than instruction finetuning on native script data, highlighting efficiency and better cross-lingual transfer. Figure 1 summarizes these results of various models on romanized and native inputs.
  • Specifically, for generation tasks we see significant improvement due to utilization of romanization. To the best of our knowledge, ours is the first work that shows that romanization can help natural language generation tasks.
  • Romanized representation can enable cross-lingual transfer in decoder-only English-heavy LLMs. Previous work has primarily focused on use of romanization in multilingual, encoder-only models.
We will release our code, datasets and model checkpoints upon the paper’s acceptance.
2 Related Work
Transliteration refers to the conversion of text written in one script to another. Romanization is a specific instance of transliteration, where the target script is the Roman/Latin script. Romanization has special significance since Roman script is by far the most widely adopted script in the world and many language models are primarily trained for English, which is written in the Roman script. Transliteration is typically used to represent different languages in the same script to enable cross-lingual transfer.
In NMT, Amrhein and Sennrich (2020) show that transliteration shows improvements for low-resource languages with different scripts by transferring from related high resource languages that use different scripts. Goyal et al. (2020) show that transliteration helps even when only contact relationship exists between the languages involved. Song et al. (2020) show that transliteration during pre-training stage for NMT also helps cross-lingual transfer.
Transliteration has also been used for cross-lingual transfer in the context of pre-trained language models. Some works (Khemchandani et al., 2021; Dhamecha et al., 2021; Moosa et al., 2022; Purkayastha et al., 2023) employ transliteration to a common script during the pretraining phase to enable cross-lingual transfer. Other studies (Dabre et al., 2022; Muller et al., 2021; Chau and Smith, 2021) adopt transliteration during the fine-tuning. Transliteration could be done to a common non-Latin script (Khemchandani et al., 2021; Dhamecha et al., 2021; Dabre et al., 2022; Doddapaneni et al., 2023) or to the Latin script Muller et al. (2021); Moosa et al. (2022); Purkayastha et al. (2023).
While transliteration has been explored for language modeling as described above, our work differs from previous work in some important aspects:
  • Previous work explored cross-lingual transfer using transliteration with multilingual language models. We focus our attention on English LLMs, and try to achieve cross-lingual transfer via romanization using English LLMs. This is a challenging scenario, since the language can be unrelated to English, and very little (if any) native or romanized data in the language under study might be seen during pre-training of the English LLM. At the same time, this is a very practical need since most best performing LLMs are English-heavy Touvron et al. (2023); Jiang et al. (2023) and hence cross-lingual transfer via romanization is an important direction to explore.
  • We investigate the utility of transliteration in decoder-only language models, which is currently the standard architecture for LLMs. In contrast, all previous work explored cross-lingual transfer with transliteration in the context of encoder-only models (the exception is Dabre et al. (2022) which use encoder-decoder models).
  • While previous work has explored the use of romanization for cross-lingual transfer in language understanding tasks, we show that romanization can be beneficial for generation tasks as well.
To the best of our knowledge, no prior research has investigated leveraging of romanization for mostly-English LLMs in the context of cross-lingual transfer to non-English languages.
Figure 2: Overview of our proposed approach.
3 Utilizing Romanized Data to make LLMs Multilingual
Our proposed approach aims to enhance the capabilities of mostly-English LLMs for non-English languages by leveraging romanized data. We first continually pretrain the LLM with both romanized and English data to create a base LM that is romanization-aware. Subsequently, we extend the process by instruction tuning the continually pretrained model. The framework of our approach is illustrated in Figure 2, encompassing the stages of data romanization, pre-training, and instruction tuning.
3.1 Romanization Scheme
A variety of romanization schemes are available, each driven by multiple considerations. One key factor is the resemblance of the romanized representation to the way people typically write romanized text. This is particularly advantageous if the pre-training data includes romanized text, as it might aid in aligning with English. Another important aspect is the fertility achieved by the romanization scheme when considering the original LLM’s tokenizer. A third consideration is whether the transliteration scheme is lossy or lossless. A lossless scheme is preferable when the objective is to convert the output back to the native script. Typically, deterministic transliteration schemes are lossless, whereas natural transliteration schemes tend to be lossy.
In this work, we have focused on Indic languages. Hence, we evaluated two romanization schemes for Indic languages: (a) the extended ITRANS scheme from the IndicNLP library Kunchukuttan (2020), which defines a fixed, reversible mapping between Indic scripts and Roman characters, and (b) the IndicXlit scheme Madhani et al. (2023), which generates romanizations as commonly used by Hindi speakers in informal contexts and is learned from parallel transliteration corpora. These mappings are inherently lossy. The IndicXlit romanization demonstrates lower fertility compared to ITRANS romanization (See Table 1). Preliminary prompting experiments with the Llama 2 model also indicated that IndicXlit outperforms ITRANS in romanized Hindi to English translation. Consequently, we selected IndicXlit as our romanization scheme for this project. However, it is important to note that the transliterations are not reversible. With continued pre-training, ITRANS transliterations might eventually achieve similar task performance to IndicXlit while ensuring script reversibility. This direction might also be relevant for other languages which do not have a transliteration model available, while romanizations mappings are available for almost all scripts. We leave this exploration for future work.
3.2 Continual Pretraining
While the base model is trained on roman script, most of the data it has been exposed to is in English. Hence, the base model cannot be used as-is for processing romanized inputs, particularly for generation into non-English languages. Hence, the base model needs to be updated. To make the base model romanization-aware, we continue to pre-train the base LLM with romanized document-level data. To prevent any catastrophic forgetting of English capabilities, we also incorporate an equal amount of English data into the pre-training mix. We perform full-finetuning of the models. We do not expand or change the model vocabulary, hence the romanized inputs can take advantage of the token embeddings from the base model while adapting to the new languages.
3.3 Instruction Finetuning (IFT)
Non-English languages possess very limited finetuning data. A common method to overcome this limitation is to translate English IFT datasets into the languages. We could choose to translate both input and outputs Wei et al. (2023), or translate only the inputs and ask a powerful, proprietary LLM like ChatGPT to generate translations in the native language Li et al. (2023). We chose the former approach since the quality of LLM responses (even for the best commerical models) might not be of good quality Ahuja et al. (2023), while high-quality open-source translation models Costa-jussà et al. (2022); Gala et al. (2023) are available making translation of both inputs and outputs feasible at scale for many languages. The translated IFT datasets are then romanized using IndicXlit.
4 Experimental Settings
We conducted comprehensive experiments spanning multiple languages, benchmarks and models. This section elaborates on the experimental settings. In all our experiments, we experiment with Llama 2 7B (Touvron et al., 2023) as the base model under various settings.
4.1 Languages
We experimented with the following Indic languages: Hindi, Marathi, Gujarati, Tamil and Malayalam. These languages span 2 languages families and 4 different scripts and are mid to low-resource languages. Hence these are representative of various multilingual scenarios where romanization can be put to use.
4.2 Datasets
Continual Pretraining: For continual pretraining, we sourced approximately 500 million words of document-level data from web-crawled corpora (Doddapaneni et al., 2023) for each language under consideration along with English. To generate the romanized dataset, we transliterated the native script dataset using the IndicXlit model (Madhani et al., 2023), a state-of-the-art open-source transliteration model for Indian languages. Both the native script dataset and its romanized counterpart were then used for continual pretraining in various configurations explained later.
Instruction Fine-tuning: These languages have very little native instruction tuning data for diverse tasks. Following Wei et al. (2023), we rely on translating high-quality English-supervised instruction-tuning datasets into the langauges under consideration. We use IndicTrans2 (Gala et al., 2023), the state-of-the-art open-source MT model for Indian languages compared to commercial offerings for translation. We sampled examples from various English instruction tuning datasets to ensure a diverse mix of tasks from the Flan collection (Longpre et al., 2023) (65k) and Dolly (Conover et al., 2023) (15k). These are translated into all the languages. The instruction tuning dataset includes a translation subset as well to learn the translation task as well as to drive alignment between English and other languages. We sample 20k high-quality manually translated examples in each direction from the BPCC-Human subset (Gala et al., 2023). Thus, the final IFT dataset has 120k examples per language. The instruction tuning datasets are further romanized using IndicXlit.
Evaluation Data: We evaluate our models on a variety of NLU and NLG tasks. These include native language test sets from the IndicXTREME (Doddapaneni et al., 2023) and Indic NLG Suite (Kumar et al., 2022) benchmarks. For the translation task, we use the FLORES-200 devtest (Goyal et al., 2022a; Costa-jussà et al., 2022) for evaluation. Following recent work (Lai et al., 2023; Üstün et al., 2024; OpenAI et al., 2023), we evaluate the knowledge and reasoning capabilities of the model on the machine-translated versions of the benchmarks such as MMLU (Hendrycks et al., 2021a), ARC (Clark et al., 2018a), BoolQ (Clark et al., 2019a) and CommonSenseQA (Talmor et al., 2019). Note that as the CommonSenseQA benchmark comprises of a blind test set, we report the evaluations on the dev set. The translations were obtained using IndicTrans2 and the romanized inputs are created using IndicXlit.
4.3 Training and Finetuning Details
The models are fully-finetuned during continual pre-training as well supervised finetuning. Both the continued pre-training and instruction finetuning is multilingual i.e. a single model is trained jointly with data for all languages (including English to preserve English performance). Note that no vocabulary expansion is done. For pre-training, loss is computed for all the tokens in the sequence. During instruction fine-tuning, the loss was computed only for the output tokens. Detailed information about the hyperparameters used for both continual pretraining and instruction fine-tuning can be found in Appendix A. We adapt open-instruct [^2] https://github.com/allenai/open-instruct/tree/main (Wang et al., 2023) for our continual pre-training and finetuning experiments.
4.4 Trained Models
We begin with Llama 2 as the base model. Then, we create two variants that have undergone continual pre-training: one with native script data (CPT-N) and the other with romanized data (CPT-R). For each of these continually pre-trained models, we further perform instruction fine-tuning (IFT). We fine-tune CPT-N using native script IFT data to obtain IFT-N, and CPT-R is fine-tuned with romanized script IFT data to produce IFT-R.
4.5 Evaluation Metrics
The evaluation metric for machine translation in our experiments is chrF (Papineni et al., 2002) computed using the SacreBLEU toolkit (Post, 2018).333nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|
version:2.3.1
For other NLG tasks, we report Rouge-L Lin (2004) scores. We primarily report evaluations on the native script outputs. In the case of romanized outputs, we transliterate them back to native script using IndicXlit combined with a unigram word-level language model Madhani et al. (2023). Combining word-level results with a unigram language model is efficient and preserves most gains of incorporating a complete n-gram language model Kirov et al. (2024). In this native script evaluation, the roman script models outputs can also have transliteration errors, which put them at a disadvantage compared to native script model. To purely compare task performance, we also report results on romanized outputs. For this, we romanize the native script outputs using IndicXlit. Since the final output is an artifact of the same romanization model, this evaluation compares only the task performance.
4.6 Models Compared
We use the 7B-parameter Llama 2 base model as our starting point for all experiments. We evaluate the following models: First, the base Llama 2 model (base) is assessed using both native script and romanized inputs. We then evaluate the continually pre-trained models with their respective script data; thus, CPT-N is evaluated with native script data, and CPT-R is evaluated with romanized data. Additionally, the instruction-tuned models are evaluated with their respective script data; therefore, IFT-N is assessed with native script data, and IFT-R is assessed with romanized data.
4.7 Prompting and Decoding
For the base and CPT models, we do 1-shot and 3-shot prompting. For IFT models, we experiment with 0-shot prompting. The prompt templates for various tasks are shown in Appendix B. For generation tasks, we use greedy decoding. For classification tasks, we completed the prompts with various possible options as per class labels and compute the output probabilities. We choose the class that maximizes the probability (Liu et al., 2021).
5 Results
In this section, we present results of various experiments comparing romanized and native representations and discuss the results.
We aim to address the following research questions:
  • Is there an efficiency gain through romanization, and if so, which approach to romanization is the most efficient?
  • Do romanized inputs lead to embeddings that are better aligned to English representations compared to native inputs?
  • Does romanization improve downstream task performance across multiple NLP tasks?
Language N R R-IndicNLP
Gujarati 18.44 3.39 4.16
Hindi 7.36 2.98 3.98
Malayalam 12.85 5.04 5.56
Marathi 8.91 3.64 4.84
Tamil 12.11 4.89 5.35
Table 1: Fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2022) for native (N), romanized with IndicXLIT (R) and romanized with IndicNLP library (R-IndicNLP). Bold indicates lowest fertility.
Efficiency Gains with Romanization
To address this question, we compute fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2022). Fertility refers to the number of tokens per word generated by the Llama 2 tokenizer Touvron et al. (2023). We compare the native script against two forms of romanized text: transliteration using IndicXlit Madhani et al. (2023) and rule-based extended ITRANS transliteration scheme implemented in the IndicNLP library Kunchukuttan (2020). Table 1 presents the fertility scores on the Llama 2 model for the languages we experimented with. We observed that the fertility (and hence sequence length) in the native script is at least twice that of the IndicXlit romanized text. In languages like Gujarati, the native script feritility is very large (greater than 4 times the romanized fertility). Hence, we can expect significant efficiency gains when processing romanized text over native text in terms of reduced memory consumption, reduced generation time and increased maximum sequence length limit.
Furthermore, romanization using IndicNLP Kunchukuttan (2020) has slightly higher fertility compared to IndicXlit. While we have used IndicXlit-based transliterations in this work for reasons explained earlier, it would be interesting to explore deterministic and reversible romanizations like IndicNLP at the cost of slightly increased inefficiency. The advantage of reversible mapping would be avoiding transliteration errors when mapping the outputs back to native script.
Language E - N E - R N - R
Gujarati 0.39 0.47 0.51
Hindi 0.40 0.50 0.34
Malayalam 0.40 0.46 0.52
Marathi 0.44 0.48 0.58
Tamil 0.44 0.43 0.53
Table 2: Cosine similarities are computed using the last token representations of the last layer from Llama 2. In this context, E - N represents the cosine similarity between English and the native script, E - R represents the similarity between English and the romanized script and N - R represents the similarity between the native and the romanized script.
Cross-lingual Representation Similarity
We compare the sentence embeddings generated by the native and romanized inputs as well as English on the base model. Specifically, we compare cosine similarities between sentence embeddings generated by these inputs. Sentence embeddings are computed using the last token representation of the final hidden layer. In Table 2, we see that the Roman script representations are closer to English representations compared to the native script representations, making them a better candidate for cross-lingual transfer from English.
Machine Translation
We discuss results of machine translation, a canonical multilingual language generation task. Table 3 (Rows XX-En translation and En-XX translation) shows results on machine translation both in and out of English.
For translation into English, the base model translation quality is roughly equivalent for native and Roman scripts. We see that continual pre-training helps improve translation quality significantly for both representations. Romanized scripts show significantly more improvement on the CPT models compared to native scripts. We hypothesize that romanization is better able to align cross-lingual representations between English and other languages to improve translation performance. Romanized IFT models also perform better than their native counterparts, albeit the gap is now narrower compared to CPT models.
For translation out of English, the base model performs very poorly which is understandable since the model might have seen only negligible non-English data. CPT significantly improves translation quality on both representations, with romanized representation showing better performance. The IFT-R models are able to take good advantage of the romanized representations, and they show an average improvement of 8 chrF points over the corresponding native script IFT-N models. Thus, we show that even generation into romanized script can take advantage of the English capabilities of the LLM.
Table 4 shows the effect of the the choice of script when evaluating translation out of English to each language. We also report evaluation on Roman script outputs for both models, and see that the romanized IFT models report even larger gains in this evaluation setting. This suggests that the true gains from romanization might be larger and it is not reflected in native script evaluation on account of errors in Roman to native transliteration that is performed for the evaluation.
Script BaseLLM CPT IFT
1-shot 3-shot 1-shot 3-shot 0-shot
XX-En translation N 22.46 23.42 38.54 37.07 49.78
R 22.52 23.52 42.53 41.64 50.75
En-XX translation N 13.95 14.25 25.55 26.19 37.40
R 14.20 12.02 29.55 30.77 46.87
XLSum N 6.88 - 7.59 - 7.77
R 10.16 - 12.44 - 12.56
IndicHeadline N 13.66 - 18.04 - 12.61
R 15.56 - 18.92 - 16.03
Table 3: Results on NLG Benchmarks - averaged across languages (RougeL score for XLSum and IndicHeadline; chrF for MT. En-$X$ refers to out-of-English and $X$-En refers to into-English directions. N and R indicate whether the model was trained and decoded on native or romanized script. Evaluation is always in native script.
Language Native-Script Eval Roman-Script Eval
N R N R
Gujarati 23.9 26.0 45.4 53.5
Hindi 44.1 49.1 45.9 56.2
Malayalam 39.4 42.8 49.0 53.8
Marathi 40.3 44.8 44.0 50.7
Tamil 39.3 41.5 50.1 53.7
Table 4: Native and Roman script evaluation results via chrF scores for En-XX translation on IFT model in 0-shot setting. N and R indicate whether the model was trained and decoded on native or romanized script.
Language Generation Tasks
Table 3 shows the results on other generation tasks viz. summarization and headline generation. Similar to MT, we observe that models utilizing romanized script consistently outperform those using native script by ~4 RougeL points across both continually pretraining and instruction fine-tuning setups with different shots. On summarization using the XLSum dataset (Hasan et al., 2021), we find that CPT and IFT models demonstrate competitive performance in both native and romanized scripts. A bit surprisingly, we find that IFT models underperform by 3-5 RougeL points compared to CPT models on IndicHeadline. We hypothesize that the format of the example was not seen in the IFT dataset, resulting in poorer zero-shot performance compared to the one-shot performance of the CPT model. Overall, we find romanization to benefit not only in terms of efficiency but also in terms of downstream performance.
Script BaseLLM CPT IFT
1-shot 3-shot 1-shot 3-shot 0-shot
IndicSentiment N 45 44.32 70.58 87.32 88.88
R 48.16 43.7 88.46 92.82 82.44
IndicCOPA N 21.08 25.43 20.47 35 37.3
R 17.53 8.12 18.23 18.78 45.45
IndicXNLI N 16.98 21.1 18.68 35.76 42.3
R 27.24 31.16 36.8 37.28 38.38
[…]
Script BaseLLM CPT IFT
1-shot 3-shot 1-shot 3-shot 0-shot
MMLU N 26.87 27.43 28.70 28.98 24.59
R 26.94 27.48 30.59 31.14 26.28
BoolQ N 57.90 44.45 59.09 47.48 60.04
R 60.75 61.03 61.85 57.10 46.95
Arc Easy N 25.82 26.35 28.77 28.25 26.01
R 26.38 26.44 26.69 27.10 25.34
Arc Challenge N 26.54 25.70 27.56 27.20 24.20
R 24.61 25.56 26.40 26.66 22.80
CommonSenseQA N 21.01 20.45 26.55 27.42 24.16
R 20.07 20.18 25.55 28.18 25.95
[…]
Language Understanding Tasks
Table 5 shows the results on different NLU tasks part of IndicXTREME benchmark Doddapaneni et al. (2023), including IndicSentiment, IndicCOPA, and IndicXNLI. We find that romanized script models are competitive to those that utilize native script. Specifically, in tasks such as IndicSentiment and IndicXNLI, romanized models consistently outperform their native counterparts across various shots. However, native models generally perform superior compared to romanized variants for the IndicCOPA task, except for the IFT model. Furthermore, we observe notable performance improvements for IndicQA without context (a reading comprehension task) across all models with romanized script when we provide the relevant passage in the context. This indicates that models with the romanized script tend to benefit from knowledge transfer from English and are also efficient in terms of sequence length.
Knowledge and Reasoning tasks
Table 6 shows the results on translated versions of various knowledge and reasoning benchmarks such as MMLU (Hendrycks et al., 2021b), BoolQ (Clark et al., 2019b), ARC (Clark et al., 2018b) and CommonsenseQA (Talmor et al., 2019). For most of the tasks, the results of romanized script are competitive with native script. At the same time, they are efficient to process, as discussed earlier. For IndicQA without context (open domain question answering task), we observe that the romanized model shows ~10 F1 point gain over the native script model. This demonstrates the overall effectiveness of utilizing romanization for non-Latin script languages.
Discussion
From the analysis and results presented earlier, we can summarize the following. Romanized representations are significantly more efficient compared to native script representations. They are better aligned to English representations compared to native script representations, hence they are better for cross-lingual transfer. Across a range of tasks and languages, romanized representations provide competitive or better performance over their native script counterparts (detailed results per language and task are presented in Appendix D. Specifically, for a generation task like machine translation we see significant gains while translating out of English with romanized representations. In Appendix C, we show some examples of romanized and native script translations, illustrating the gains from romanization. These results indicate that romanization is a promising alternative to extending the capabilities of English LLMs to other languages using non-Roman scripts.
6 Conclusion and Future Work
In this study, we proposed the use of romanization to extend the performance of LLMs primarily trained in English to other languages. Our approach successfully unlocks LLM capabilities for non-English languages by using romanization to bridge English and non-English language representations. We have empirically demonstrated the effectiveness of this strategy through experiments involving few-shot prompting, continual pretraining, and instruction fine-tuning on a variety of tasks. Additionally, using romanized data increases inference speed and maximum processable sequence length, while reducing memory requirements by 2x to 4x depending on the language.
Looking forward, we also plan to explore reversible, deterministic transliteration so that transliteration errors in output post-processing are eliminated. In addition, it would be interesting to explore if romanized representation can act as a bridge to efficiently improve native script performance also when English LLMs are extended to incorporate non-Roman tokens.
7 Limitations
In this work, we chose to use natural romanizations since they tended to have lower fertility and might have been seen in the base LM pre-training. Transliterating the outputs back to native script can be lossy. We use a high-quality transliteration model to minimize the transliteration errors. However, such high quality transliterations might not be available for all languages. Using a deterministic and reversible transliteration scheme can alleviate this problem.
The experiments in this work have been conducted on Indian languages. Indian languages are representative of the scenarios we study and we cover multiple language families and script with varying resource status. Previous studies have also used Indic languages as a case study for multilinguality and multi-script scenarios Moosa et al. (2022); Khemchandani et al. (2021). Hence, we believe the findings should generalize to other language families - but further experimentation can confirm this hypothesis.
We have done only a limited amount of pretraining data in our experiments due to compute constraints. Resource constraints led us to limit our experiments to a 7B model, but experiments with larget models can yield more insights. Larger models and pre-training on larger datasets might benefit the multilingual models Kaplan et al. (2020); Fernandes et al. (2023). We expect the broad conclusions to be the same - that romanized models are more efficient and better/comparable in task performance to native-script models.
Some of the evaluations have been done on translated benchmarks. While this is not ideal, most low-resource languages testsets across diverse tasks. Hence, multiple works have relied on translated benchmarks to measure multilingual performance - these can give indicative trends on performance across languages. More efforts are needed to created multilingual benchmarks for diverse tasks.
8 Ethics Statement
We use romanization as a way to align English with other languages using non-Latin scripts. Given the current state of LLMs, this seems like a practical direction to extend capabilities of the best LLMs to other languages. The intention is not to supplant the use of a native script (which is widely adopted and has a rich literary tradition) with romanized script, but to use romanization as a way to efficiently bring the benefits of LLM technology to low-resource languages written in non-Latin scripts. Further advancements are needed to extend this line of research to improve native script performance efficiently.
This work does not involve any new data collection and does not employ any annotators for data collection. We utilize publicly available datasets for the experiments reported in this work. Some of these datasets originate from web crawls, and we do not explicitly attempt to identify any biases within these datasets, using them in their original form.
Acknowledgments
We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research and building datasets, models, tools and other resources for Indian languages.
Appendix A Model Training Details
Tables 7 and 8 show the hyperparameters used for training the CPT and IFT models.
Hyperparameter Value
Batch Size (tokens) 1M
Learning Rate 5e-5
Number of Epochs 1
Maximum Sequence Length 2,048
Table 7: The range of hyperparameters used for continual pretraining.
Hyperparameter Value
Batch Size (examples) 128
Learning Rate 5e-5
Number of Epochs 1
Maximum Sequence Length 2,048444Although this is the maximum permissible length, most examples fall far below this length during fine-tuning.
Table 8: The range of hyperparameters used for supervised fine-tuning.
Appendix B Prompt Templates
We list the various prompt templates for evaluations across different test sets (see Figures 3, 4, 5, 6, 7, 8, 9, 10, 11 and 12).
{promptbox}
Predict the sentiment of the review. The possible choices for the sentiment are: ’positive’ and ’negative’.
Review: {{ text }}
Sentiment: {{ label }}
$\cdots$
$\cdots$
Review: {{ text }}
Sentiment:
Figure 3: Prompt template for IndicSentiment
{promptbox}
I am hesitating between two options. Help me choose the more likely cause or effect.
{{ premise }} {{ connector }}
A. {{ choice1 }}
B. {{ choice2 }}
Answer: {{ label }}
$\cdots$
$\cdots$
{{ premise }} {{ connector }}
A. {{ choice1 }}
B. {{ choice2 }}
Answer:
Figure 4: Prompt template for IndicCOPA
{promptbox}
Answer whether the hypothesis is more likely to be true (entailment), false (contradiction), or unknown (neutral) based on the given premise.
Premise: {{ premise }}
Hypothesis: {{ hypothesis }}
Answer: {{ label }}
$\cdots$
$\cdots$
Premise: {{ premise }}
Hypothesis: {{ hypothesis }}
Answer:
Figure 5: Prompt template for IndicXNLI
{promptbox}
Answer the following question based on the information in the given passage.
Passage: {{ passage }}
Question: {{ question }}
Answer: {{ answer }}
$\cdots$
$\cdots$
Passage: {{ passage }}
Question: {{ question }}
Answer:
Figure 6: Prompt template for IndicQA.
{promptbox}
Summarize the following {{ language }} article(s) as accurately as possible in few sentences.
{{ language }} article: {{ article }}
{{ language }} summary: {{ summary }}
$\cdots$
$\cdots$
{{ language }} article: {{ article }}
{{ language }} summary:
Figure 7: Prompt template for XLSum.
{promptbox}
Generate a headline for the following article(s) as accurately as possible.
{{ language }} article: {{ article }}
{{ language }} headline: {{ headline }}
$\cdots$
$\cdots$
{{ language }} article: {{ article }}
{{ language }} headline:
Figure 8: Prompt template for IndicHeadline.
{promptbox}
The following are multiple choice questions (with answers) about {{ subject }}.
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
Answer: {{ label }}
$\cdots$
$\cdots$
{ subject }}
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
Answer:
Figure 9: Prompt template for MMLU.
{promptbox}
The following are binary yes/no choice questions (with answers).
Passage: {{ passage }}
Question: {{ question }}
Answer: {{ label }}
$\cdots$
$\cdots$
Passage: {{ passage }}
Question: {{ question }}
Answer:
Figure 10: Prompt template for BoolQ.
{promptbox}
The following are multiple choice questions (with answers) about science.
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
E. {{ choice5 }}
Answer: {{ label }}
$\cdots$
$\cdots$
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
E. {{ choice5 }}
Answer: {{ label }}
Figure 11: Prompt template for ARC.
{promptbox}
The following are multiple choice questions (with answers) requiring common sense.
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
E. {{ choice5 }}
Answer: {{ label }}
$\cdots$
$\cdots$
{{ question }}
A. {{ choice1 }}
B. {{ choice2 }}
C. {{ choice3 }}
D. {{ choice4 }}
E. {{ choice5 }}
Answer: {{ label }}
Figure 12: Prompt template for CommonsenseQA.
Appendix C Translation Examples
This section shows some examples of translations using the romanized and native-script IFT models (See Figures 13, 14, 15 and 16). We can see in these selected examples that romanized representaton performs better.
Figure 13: Example of zero-shot translation for English to Hindi machine translation task to generate text in native script using a model fine-tuned on native script data.
Figure 14: Example of zero-shot translation for English to Hindi machine translation task to generate text in romanized script using a model fine-tuned on romanized data. The romanized output is then transliterated to native script.
Figure 15: Example of zero-shot translation for the Hindi to English machine translation task using native script input for a model fine-tuned on native script data.
Figure 16: Example of zero-shot translation for the Hindi to English machine translation task using romanized script input for a model fine-tuned on romanized data.
Appendix D Language Specific Results
We report the language-wise evaluation results for 5 Indic languages considered for this study across a diverse array of tasks ranging from text generation, text understanding, and text reasoning in Tables 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21 and 22.
Model Shots Input text Hindi Gujarati Tamil Marathi Malayalam
N R N R N R N R N R
Llama 2 1 Native 22.95 29.21 8.15 11.9 12.12 14.62 15.32 20.22 11.2 14.98
Llama 2 3 Native 23.73 29.91 7.38 11.48 12.4 15.49 16.19 21.4 11.55 15.31
Llama 2 1 Romanized 17.25 18.86 14.63 16.16 12.8 14.87 13.42 15.85 12.88 15.48
Llama 2 3 Romanized 15.01 18.25 10.8 14.79 9.3 11.39 12.09 16.61 12.88 15.53
CPT - N 1 Native 32.07 37.89 15.94 18.79 25.81 28.87 29.47 34.24 24.48 28.36
CPT - N 3 Native 32.9 38.51 16.3 19.29 26.2 29.12 29.9 34.75 25.64 29.45
CPT - R 1 Romanized 30.93 33.66 29.73 35.35 30.11 32.71 28.27 33.71 28.72 32.47
CPT - R 3 Romanized 31.51 34.99 30.55 36.8 31.5 35.04 29.42 35.08 30.85 34.86
IFT (Llama 2) - N 0 Native 43.28 48.43 23.54 25.82 38.99 41.13 40.13 44.72 38.85 42.12
IFT (CPT - N) - N 0 Native 44.08 49.12 23.86 25.97 39.3 41.51 40.31 44.8 39.44 42.78
IFT (Llama 2) - R 0 Romanized 45.15 55.23 44.8 52.98 49.13 52.79 42.78 49.6 47.71 52.63
IFT (CPT - R) - R 0 Romanized 45.94 56.19 45.37 53.5 50.06 53.71 44 50.73 48.96 53.76
Table 9: chrF scores for En-XX translation task on FLORES-200 devtest (Goyal et al., 2022a) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model Shots Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 40.49 13.63 16.06 24.18 17.93
Llama 2 3 Native 40.38 17.58 17.22 24.15 17.78
Llama 2 1 Romanized 29.24 22.62 17.06 22.56 21.1
Llama 2 3 Romanized 28.91 23.28 19.86 23.6 21.96
CPT - N 1 Native 51.37 24.77 33.2 47.04 36.34
CPT - N 3 Native 50.31 25.73 28.56 43.37 37.4
CPT - R 1 Romanized 45.43 45.08 38.34 44.28 39.52
CPT - R 3 Romanized 45.4 44.82 33.72 43.56 40.71
IFT (Llama 2) - N 0 Native 51.7 47.26 42.27 48.8 47.83
IFT (CPT - N) - N 0 Native 53.67 49.51 44.44 51.29 50
IFT (Llama 2) - R 0 Romanized 51.84 50.17 41.67 48.51 47.14
IFT (CPT - R) - R 0 Romanized 53.89 53.3 45.09 51.06 50.43
Table 10: chrF scores for XX-En translation task on FLORES-200 devtest (Goyal et al., 2022a) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model Shots Input text Hindi Gujarati Tamil Marathi
N R N R N R N R
Llama 2 1 Native 13.32 9.59 3.2 2.79 4.13 3.47 6.85 6.11
Llama 2 1 Romanized 13.43 15.96 8.95 7.65 9.05 6.69 9.22 7.45
CPT - N 1 Native 16.32 12.18 5.18 4.58 4.3 3.91 4.55 4.61
CPT - R 1 Romanized 13.05 14.77 13.8 12.98 10.86 9.68 12.06 11.34
IFT (Llama 2) - N 0 Native 11.71 8.66 4.38 3.18 6.62 5.62 7.41 6.61
IFT (CPT - N) - N 0 Native 12.56 9.21 4.3 3.52 6.51 5.66 7.71 6.64
IFT (Llama 2) - R 0 Romanized 14.73 16.87 11.84 10.56 10.83 8.59 11 9.53
IFT (CPT - R) - R 0 Romanized 15.05 17.92 12.36 11.16 11.28 9.32 11.56 9.87
Table 11: RougeL scores on XLSum (Hasan et al., 2021) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
N R N R N R N R N R
Llama 2 1 Native 20.91 15.01 6.17 5.55 19.14 17.8 9.26 8.45 12.84 11.45
Llama 2 1 Romanized 17.58 20.55 12.18 13.72 23.5 26.16 10.55 11.32 13.97 18.83
CPT - N 1 Native 22.74 16.52 9.07 7.97 23.59 22.79 19.21 18.72 15.57 14.74
CPT - R 1 Romanized 20.19 24.44 12.83 14.34 29.13 32.03 15.14 17.39 17.29 22.38
IFT (Llama 2) - N 0 Native 18.38 13.43 6.36 5.64 15.88 15.35 9.96 9.84 10.51 9.85
IFT (CPT - N) - N 0 Native 19.47 14.24 6.65 5.98 15.14 14.75 10.77 10.7 11.04 10.68
IFT (Llama 2) - R 0 Romanized 17.76 21.46 12.88 15.43 20.89 22.27 11.51 12.62 13.45 17.45
IFT (CPT - R) - R 0 Romanized 18.74 21.91 13.04 15.64 21.61 23.82 12.2 13.59 14.56 19.27
Table 12: RougeL scores on IndicHeadline (Kumar et al., 2022) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 28.67 25.59 26.06 27.58 26.44
Llama 2 3 Native 29.11 26.57 26.75 27.85 26.86
Llama 2 1 Romanized 27.52 26.57 26.34 27.14 27.11
Llama 2 3 Romanized 28.38 27.62 26.65 27.24 27.52
CPT - N 1 Native 29.61 27.34 28.63 29.58 28.35
CPT - N 3 Native 30.08 27.62 28.66 29.91 28.61
CPT - R 1 Romanized 30.8 30.72 30.38 31.33 29.7
CPT - R 3 Romanized 31.55 31.05 30.18 32.15 30.76
IFT (Llama 2) - N 0 Native 24.47 24.08 24.09 23.3 24.36
IFT (CPT - N) - N 0 Native 24.72 24.05 24.78 24.62 24.81
IFT (Llama 2) - R 0 Romanized 24.75 25.96 26.03 25.54 26.48
IFT (CPT - R) - R 0 Romanized 25.58 26.4 26.73 26.1 26.57
Table 13: Accuracy on translated MMLU (Hendrycks et al., 2021a) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 57.58 58.78 56.88 61.07 55.2
Llama 2 3 Native 45.08 57.58 39.3 40.92 39.39
Llama 2 1 Romanized 62.23 60.03 58.72 60.61 62.14
Llama 2 3 Romanized 61.93 61.77 60.09 58.9 62.48
CPT - N 1 Native 60.7 56.21 57.71 61.41 59.45
CPT - N 3 Native 50.67 60.64 45.78 39.36 40.95
CPT - R 1 Romanized 62.17 61.5 61.44 61.93 62.23
CPT - R 3 Romanized 64.77 60.92 55.41 47.86 56.54
IFT (Llama 2) - N 0 Native 60.12 61.65 60.06 61.01 60.18
IFT (CPT - N) - N 0 Native 60.34 61.5 58.32 60.43 59.6
IFT (Llama 2) - R 0 Romanized 61.62 56.91 58.01 58.69 47.68
IFT (CPT - R) - R 0 Romanized 60.7 45.02 38.01 53.24 37.77
Table 14: Accuracy on translated BoolQ (Clark et al., 2019a) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 27.27 25.67 25.63 25.13 25.42
Llama 2 3 Native 27.36 26.09 26.18 26.6 25.51
Llama 2 1 Romanized 27.23 25.29 26.05 26.98 26.35
Llama 2 3 Romanized 27.27 26.64 25.29 26.85 26.14
CPT - N 1 Native 30.98 26.22 28.87 30.56 27.23
CPT - N 3 Native 30.68 26.43 27.4 29.12 27.61
CPT - R 1 Romanized 27.23 25.88 26.26 26.89 27.19
CPT - R 3 Romanized 27.61 26.56 26.39 27.74 27.23
IFT (Llama 2) - N 0 Native 25.42 25.88 25.25 25.42 26.09
IFT (CPT - N) - N 0 Native 26.35 25.84 25.67 25.93 26.26
IFT (Llama 2) - R 0 Romanized 25.08 25.72 25.04 25.34 25.76
IFT (CPT - R) - R 0 Romanized 25.97 25.34 24.24 25.42 25.72
Table 15: Accuracy on translated ARC-Easy (Clark et al., 2018b) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 26.45 27.05 26.71 26.71 25.77
Llama 2 3 Native 27.22 26.19 25.17 25.6 24.32
Llama 2 1 Romanized 24.4 25.68 24.32 24.32 24.32
Llama 2 3 Romanized 24.83 27.47 25.09 25.09 25.34
CPT - N 1 Native 27.65 26.62 27.3 27.3 28.92
CPT - N 3 Native 29.69 25.85 26.96 26.96 26.54
CPT - R 1 Romanized 25.34 26.79 24.57 28.07 27.22
CPT - R 3 Romanized 24.49 27.56 25.94 27.99 27.3
IFT (Llama 2) - N 0 Native 22.27 22.78 23.46 21.93 23.29
IFT (CPT - N) - N 0 Native 25.85 23.21 24.57 24.06 23.29
IFT (Llama 2) - R 0 Romanized 21.67 22.87 23.04 22.87 21.67
IFT (CPT - R) - R 0 Romanized 22.78 22.95 22.7 22.35 23.21
Table 16: Accuracy on translated ARC-Challenge (Clark et al., 2018b) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 23.99 19 20.55 21.45 20.06
Llama 2 3 Native 23.75 18.75 20.22 20.22 19.32
Llama 2 1 Romanized 21.54 20.48 19.49 19.74 19.08
Llama 2 3 Romanized 21.54 19.57 19.98 19.82 19.98
CPT - N 1 Native 28.09 23.75 28.5 26.94 25.47
CPT - N 3 Native 31.12 22.44 28.99 28.09 26.45
CPT - R 1 Romanized 23.75 27.19 25.3 25.22 26.28
CPT - R 3 Romanized 25.71 30.22 28 27.1 29.89
IFT (Llama 2) - N 0 Native 20.64 19.25 20.07 20.97 20.31
IFT (CPT - N) - N 0 Native 23.34 23.18 23.59 25.14 25.55
IFT (Llama 2) - R 0 Romanized 20.23 21.54 21.29 22.93 22.44
IFT (CPT - R) - R 0 Romanized 23.91 26.7 25.55 26.04 27.52
Table 17: Accuracy on translated CommonsenseQA (Talmor et al., 2019) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 53.2 35.1 46 50.3 40.4
Llama 2 3 Native 77.8 28.4 34.5 49.2 31.7
Llama 2 1 Romanized 59.1 44.9 42.4 53.5 40.9
Llama 2 3 Romanized 67.5 38.9 33.8 44.3 34
CPT - N 1 Native 64.9 68 77.8 75.1 67.1
CPT - N 3 Native 89.7 86.6 87.8 89.6 82.9
CPT - R 1 Romanized 91.2 86.6 88.4 90 86.1
CPT - R 3 Romanized 91.9 93.3 93.3 93.1 92.5
IFT (Llama 2) - N 0 Native 90.4 88.1 87.6 89.5 89
IFT (CPT - N) - N 0 Native 87.2 88.7 89.6 90 88.9
IFT (Llama 2) - R 0 Romanized 75.3 68.8 71.2 77.6 77.3
IFT (CPT - R) - R 0 Romanized 83.7 73.5 87.7 81.9 85.4
Table 18: F1 score on IndicSentiment (Doddapaneni et al., 2023) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 41.1 41.1 1.6 41.1 1.6
Llama 2 3 Native 52.1 56.7 7.5 34.7 1.6
Llama 2 1 Romanized 39.2 36.2 1.6 26.6 1.6
Llama 2 3 Romanized 23.6 16.1 3.1 5.9 0
CPT - N 1 Native 40.5 40.7 1.6 38.4 1.6
CPT - N 3 Native 61 65.7 11.6 60 11.7
CPT - R 1 Romanized 38.6 37.4 0 31.8 1.6
CPT - R 3 Romanized 30.3 42.7 6.6 32.3 0.8
IFT (Llama 2) - N 0 Native 5.2 1.7 44.1 1.8 7.4
IFT (CPT - N) - N 0 Native 51.3 33.2 47.6 35.4 56.3
IFT (Llama 2) - R 0 Romanized 18.9 58.4 18.8 35.9 63.6
IFT (CPT - R) - R 0 Romanized 33.2 62.3 63.4 47.1 66.7
Table 19: F1 score on IndicCOPA (Doddapaneni et al., 2023) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
Llama 2 1 Native 17.6 16.7 16.8 17.1 16.7
Llama 2 3 Native 27.4 25.1 17.4 18.7 16.9
Llama 2 1 Romanized 27.5 28.5 26.5 26.4 27.3
Llama 2 3 Romanized 31.4 31.6 30.5 31.2 31.1
CPT - N 1 Native 19.9 19.3 19.4 17.7 17.1
CPT - N 3 Native 44.6 30.1 32.8 45.2 26.1
CPT - R 1 Romanized 34.1 37 36.6 37.3 39
CPT - R 3 Romanized 38.6 37.3 37.5 37.3 35.7
IFT (Llama 2) - N 0 Native 36.3 37.9 43.1 42.9 38.7
IFT (CPT - N) - N 0 Native 42.2 41.2 43.1 42.5 42.5
IFT (Llama 2) - R 0 Romanized 24 22.5 23 22.1 28.6
IFT (CPT - R) - R 0 Romanized 38.5 38.9 37.2 38.1 39.2
Table 20: F1 score on IndicXNLI (Doddapaneni et al., 2023) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
N R N R N R N R N R
Llama 2 1 Native 21.02 24.63 3.26 3.39 8.75 8.91 10.08 12.22 5.73 5.73
Llama 2 1 Romanized 29.11 32.8 19.95 21.97 11.34 13.32 21.11 21.6 17.18 20.42
CPT - N 1 Native 28.98 30.36 17.01 17.76 21.01 21.04 28.76 28.96 23.37 23.43
CPT - R 1 Romanized 29.58 32.8 27.01 28.85 24.69 27.11 30.01 30.53 24.95 27.51
IFT (Llama 2) - N 0 Native 19.75 22.79 11.66 14.3 17.56 17.81 18.39 21.96 11.15 11.4
IFT (CPT - N) - N 0 Native 21.85 25.68 13.54 16.21 18.49 18.68 18.02 21.43 12.07 12.49
IFT (Llama 2) - R 0 Romanized 23.54 33.79 19.15 24.3 16.44 25.54 21.26 25.79 13.2 20.31
IFT (CPT - R) - R 0 Romanized 24.44 33.81 21.41 26.9 17.65 25.27 23.64 28.41 14.53 22.01
Table 21: F1 score on IndicQA (with context) (Doddapaneni et al., 2023) across 5 Indic languages in both native (N) and romanized (R) script with different shots.
Model n-shot Input text Hindi Gujarati Tamil Marathi Malayalam
N R N R N R N R N R
Llama 2 1 Native 4.97 6.69 0.2 0.22 1.02 1.09 0.9 1.16 0.31 0.36
Llama 2 1 Romanized 6.96 7.33 16.77 16.82 12.26 12.26 16.45 16.38 12.34 12.77
CPT - N 1 Native 7.44 8.67 2.73 3.15 4.66 4.64 7.41 7.82 5.55 5.67
CPT - R 1 Romanized 13.85 15.19 13.46 13.41 16.66 17.13 26.75 26.82 17.38 17.93
IFT (Llama 2) - N 0 Native 2.36 2.64 0.66 0.72 1.48 1.53 1.18 1.28 0.89 0.9
IFT (CPT - N) - N 0 Native 2.73 3.03 0.85 0.95 1.15 1.16 1.27 1.47 0.82 0.9
IFT (Llama 2) - R 0 Romanized 2.21 2.93 0.65 0.76 1.37 1.42 1.75 1.75 1.01 1.33
IFT (CPT - R) - R 0 Romanized 3.32 3.69 1.03 0.96 1.4 1.67 1.81 1.9 1.05 1.21
Table 22: F1 score on IndicQA (without context) (Doddapaneni et al., 2023) across 5 Indic languages in both native (N) and romanized (R) script with different shots.